# Rocky planet formation in compact disks around M dwarfs

### v2

@@ 2.2 Dust disk model
with  the headwind pre-factor that measures the disk pressure gradient (Nakagawa et al., 1986) with  and  the gradients of the gas disk surface density and temperature and  the gas aspect ratio. In order to take into account the ice sublimation of pebbles when they cross the snowline , we multiply  by a factor of 0.5 when . Moreover, we consider that pebbles inside the snowline are dry, while outside the snowline have 50 of water in mass.
The  is defined as:
@@
We assume an initial time for our disk model of 1 Myr. We follow the gas disk model proposed by Ida et al. (2016) based on two dominant heating mechanisms: viscous dissipation for the inner disk and stellar irradiation for the outer disk (see Appendix A for a detail description of the model and the parametrization of , and ).
@@
Unlike Lambrechts & Johansen (2014), who proposed a pebble flux that varies in time depending on the location of the pebble front (radial distance where the pebbles are originated), we assume a constant pebble flux in time, that varies with the stellar mass. For a star of 0.1 M⊙ we set a pebble flux of M⊕ yr-1, while for a star of 0.3 M⊙ and a star of 0.5 M⊙ we set a pebble flux a factor 6 and a factor 10 higher, respectively, motivated by the fact that the pebble flux scales with the gas density, as suggested in Lambrechts & Johansen (2014). This scale is also consistent with the dust mass and stellar mass relations found by Ansdell et al. (2017). The previous section has demonstrated that these values in compact dust disks correspond to a millimeter dust flux similar to observed values in the unresolved compact disks in Lupus. We note that such values are also in agreement with the order of magnitude of the pebble fluxes estimated by Mulders et al. (2021) for low-mass stars at 1 Myr, and the value proposed recently by Pan et al. (2024) that studied planet formation around 0.1 M⊙ stars. This pebble flux is also consistent with ALMA observations, as demonstrated in Figure 2.
@@
Figure 3 illustrates the pebble model that we used, showing the evolution of the pebble surface density, the pebble aspect ratio and the Stokes numbers at 1 Myr, in which we assumed a star of 0.1 M⊙, a turbulent parameter  and the pebble flux of F M. The different slopes of the curves are related to the location of the snowline and to the radial distance that separates the two heating mechanisms proposed in the gas disk model (see Appendix A). We clarify that we assumed the same value for the turbulent parameter as for the gas disk viscosity, in agreement with previous works (e.g. Izidoro et al., 2017; Raymond et al., 2018; Sánchez et al., 2022).
@@
By considering the pebble fluxes mention above, we corroborated that a planet initially located at a few au reach its isolation mass in less than 200,000 yr for a star of 0.1 M⊙ (see Figure 18 as an example) and in less than 100,000 yr for a star of 0.3 M⊙ or 0.5 M⊙, as they are also exposed to a fast inward migration in a gas disk with a low viscosity (see Appendix A for a description of the gas model, Appendix B for planet-disk interactions and Appendix C for the test simulations made to validate the pebble accretion model used). We note that if we calculate the pebble mass available during this period of time from the pebble flux that we proposed, we obtain lower values than the total dust mass available in a disk of 20 au (see Section 2.1). Thus, the pebble flux value proposed is a conservative value which is lower than the maximum value needed for all the dust disk mass within 20 au to be available for the rocky planet assembly (see Section 6 for further discussion).
@@ 3 Planet formation model
low-mass planet-disk interactions, following the gas disk evolution model from Ida et al. (2016), the torques prescriptions from Paardekooper et al. (2010, 2011) and the planetary acceleration corrections prescriptions proposed by Ida et al. (2020), based on purely dynamical friction (Sánchez et al., 2022, see Appendix A and B.1 for a detailed description);
star-planet tidal interactions based on the tidal equilibrium model proposed by Hut (1981) and lately adapted by Bolmont et al. (2011) for brown dwarfs and M dwarfs, which includes the evolution of the stellar radius (Baraffe et al., 2015) and rotational period (Bolmont et al., 2011) (Sánchez et al., 2020, 2022, see Appendix B.2 for a detailed description);
@@
the evolution of the stellar luminosity and radius based on the stellar evolution models from Baraffe et al. (2015), and the evolution of the rotational period of a star of 0.1, 0.3 and 0.5 M⊙ based on the observations of rotational periods of M dwarfs listed in Bolmont et al. (2011) and in Scholz et al. (2018), for the gas disk model and the star-planet interactions (see Appendix A and B.2).
@@ 3.1 Formation scenarios
We explored different formation scenarios to study rocky planet formation. First, for our standard model we considered a star of 0.1 M⊙ and a sample of embryos with 0.01 M⊕ initially located beyond the snowline that growth by pebble accretion in a gas disk with low viscosity ( and a disk lifetime of 10 Myr (see the dust disk model in Section 2.2 and the gas model in Appendix A).
@@
as the gas and dust densities are lower in a disk with higher viscosity), to analyze the impact in the resulting planetary masses and in the migration and orbital decay of the planets. We also assumed a different initial embryo distribution to see the impact in the final fraction of water in planetary masses. Moreover, we let a sample of embryos growth by planetesimal accretion instead of pebble accretion assuming that the total mass in planetesimalsembryos is equal to the pebble mass available in the disk for accretion (see following sections). Furthermore, we considered higher stellar masses to see how rocky planet formation is scaling with the stellar mass. We considered a star of 0.3 M⊙ and a star 0.5 M⊙. In both cases we studied planet formation by pebble accretion and also by planetesimal accretion in a gas disk with low viscosity () assuming a sample of embryos initially located outside the snowline. For each scenario proposed, we run 10 N-body simulations.
@@ 3.2 N-body code characterization
To run the Mercury code, we chosed the hybrid integrator, which uses a second-order symplectic algorithm to treat interactions between objects with separations greater than 3 RHill and the Bulirsch-Stöer method to solve close encounters. We run our simulations for 50 Myr each, in order to study the dynamical stability of the system after the gas disk dissipated at 10 Myr. We adopted an integration time step of 0.1 days, in order to be able to map at least 10 times the orbit of a planet with an orbital period of 1 day. Furthermore, we consider that an embryo is ejected from the system when it reached a distance of 100 au, and we considered that an embryo had collided with the central star when it is closer than 0.005 au, which corresponds, approximately, to the maximum radius of the host star. We fixed this value for the entire simulation in order to avoid any numerical error for small-periastron orbits.
@@ 3.3 Initial embryo distribution
In the scenarios in which we considered pebble accretion and an initial embryo distribution outside the snowline, we distributed a total amount of 0.25 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.01 M⊕. The selection of initial embryos with lunar mass is in agreement with several works that consider planetary growth by pebble accretion (e.g. Liu et al., 2019; Coleman et al., 2019) and proposed initially by Kokubo & Ida (1998). Embryos are located beyond the snowline: between 0.5 and 2 au for a central star 0.1 M⊙ ( au), between 1 and 3 au for a central star of 0.3 M⊙ ( au), and between 1.5 and 3.5 au for a central star of 0.5 M⊙ ( au). In each scenario, each seed had a random separation between 5 and 15 Hill Radius, according to their classical isolation mass (e.g. Kokubo & Ida, 2000; Weiss et al., 2018). On the other hand, for the scenario in which we considered embryos located both inside and outside the snowline for a central star of 0.1 M⊙, we distributed 0.5 M⊕ in 50 seeds of protoplanetary embryos, each of them with a mass of 0.01 M⊕. In this case they are located between 0.02 and 1.5 au, randomly distributed between 10 and 15 Hill Radius.
In the scenarios in which we considered planetesimal accretion, we assumed bigger embryos close to the mass of Mars, as in previous works (e.g. Ogihara & Ida, 2009; Coleman et al., 2019). For a central star of 0.1 M⊙, we distributed a total amount of 2.5 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.1 M⊕, located beyond the snowline: between 0.5 and 2 au. For a central star of 0.3 M⊙ and a central star of 0.5 M⊙, we distributed a total amount of 5 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.2 M⊕, located between 1 and 3 au and between 1.5 and 3.5 au, respectively. In each simulation, we considered that the embryos initially located inside the snowline were dry, and the ones located outside the snowline present 50 of water in mass
In every case, we considered that all embryos initially have  and . Those values were taken from uniform distributions. The initial values for the rest of the orbital elements, argument of periastron , longitude of the ascending node , and mean anomaly , were randomly determined also from uniform distributions but between 0 and 360 degrees. The physical density of each embryo was fixed to 5  inside the snowline and 1.3  outside the snowline.
@@ 3.4.1 Pebble accretion
For the sample of planetary embryos immerse in a gas disk that growth by pebble accretion, we follow the pebble model proposed by Lambrechts & Johansen (2014). Planetary embryos with different masses accrete pebbles in different regimes. By comparing the pebble accretion radius and the pebble disk scale height , pebble accretion could be classified into two regimes: the 2D and the 3D regime. If a planetary embryo with mass Mp is large enough (), then it would accrete from the complete layer of pebbles, in the 2D regime; while if the embryo is not large enough (), it would accrete only a fraction of the pebble layer, in the 3D regime. The with the relative velocity between the pebble and the embryo.
The pebble accretion rate is defined as , where, represents the efficiency of pebble accretion (Guillot et al., 2014; Lambrechts & Johansen, 2014). Considering an embryo with an eccentric orbit and an inclination , we calculate such efficiencies following Liu & Ormel (2018) and Ormel & Liu (2018). Thus, the pebble accretion rates in the 2D regime and in the 3D regime are defined by:
@@
are the pebble efficiencies in the 2D and 3D regime, respectively, with and fitting constant of the model, , , and the relative velocity between the embryo and the pebble , with the relative velocity considering an embryo in a circular orbit and the relative velocity considering an embryo in an eccentric orbit, defined as:
with the eccentricity of the embryo’s orbit, a fitting constant of the model, the transition mass between the Bondi regime (low impact parameter between the embryo and the pebble) and the Hill regime (high impact parameter) (e.g. Ormel, 2017), relative velocity for a pebble inside Bondi regime, the relative velocity for a pebble inside the Hill regime, and , the transition velocity between the Bondi and the Hill regimes.
@@
where represents the pebble flux accreted by all the embryos exterior to the radial location of the planet. Each embryo will accrete pebbles until they reach the pebble isolation mass. This mass is considered as the necessary mass that an embryo needs to be able to perturb the local disk structure, forming a pressure bump in the disk exterior to its orbit, stopping the inward migration of pebbles, and limiting the amount of pebbles for embryos orbiting in inner orbits (Lambrechts et al., 2014). The isolation pebble mass can be defined as:
@@
As we are working with a sample of embryos, we clarify that each embryo of the simulation will accrete pebbles until it reaches the isolation mass or the total pebble flux available at the location of the planet is zero. If a planet reaches the isolation mass, then we will assume that no more pebbles could migrate inwards and thus no other inner planets could continue accreting pebbles (e.g. Lambrechts et al., 2019; Mulders et al., 2021).
@@ 3.4.2 Planetesimal accretion
In the scenarios in which we considered planetesimal accretion, we let the sample of embryos interacts with a sample of 1000 planetesimals. For a star of 0.1 M⊙, each of them has a mass of 0.01 M⊕, while for a star of 0.3 M⊙ the mass of each of them is 0.02 M⊕ and for a star of 0.5 M⊙ is 0.05 M⊕. Their semi-major axis are initially distributed following a power law  from the inner edge of the disk until the location where the total amount between embryos and planetesimals equals the pebble mass available in the disk at 1 Myr, which is between 0.02 au and 2 au for a star of 0.1 M⊙, between 0.065 au and 5 au for a star of 0.3 M⊙ and between 0.075 au and 6 au for a star of 0.5 M⊙.
In every case, the rest of the orbital elements: , , ,  and , were estimated from equivalent distributions as for the sample of embryos. The physical density of the planetesimals was fixed to 3  inside the snowline of the system and 1.5  beyond the snowline. We consider that planetesimales inside the snowline are dry, while outside the snowline present 50 of water in mass.
@@ 4 Simulated systems: Accretion history and dynamical evolution
We present the differences in the resulting simulated planetary systems of each formation scenario (see Table LABEL:tab:table2) in terms of their dynamical evolution and accretion history. Moreover, we compared their final configurations (semi-major axis and eccentricity), masses, amount of water in mass and multiplicity (see Table 3).
@@ 4.1 Standard model: PA-M01-low-icy
In our standard model we considered a central star of M M⊙, an initial embryo distribution outside the snowline that grow by pebble accretion and a gas disk with low viscosity (). In this scenario, the sample of embryos suffered an efficient inward migration in all the simulations along the disk lifetime. The resulting planets ended up located close-in to the star ( au) in less than 300,000 yr. During this period of time they suffered collisions among embryos and accreted pebbles from outside the snowline for the first 100,000 yr and then from inside the snowline (r au). Once the outermost planet reached the isolation mass, the pebbles stopped migrating inwards and, in most cases, the resulting planets achieved their final masses. The resulting planets stopped migrating inwards and reminded in compact configurations close to the inner edge of the disk, in quasi-circular and co-planar orbits. In 60 of the simulations, either 1 or 2 collisions with the central star occurred within 500,000 yr. In the rest of the simulations, no embryo collided with the central star. On the other hand, in all the simulations between the 60 and the 88 of the embryos collided among them. In 70 of the simulations, all the collisions among embryos occurred within the first 300,000 yr (while the embryos were also accreting pebbles). In the rest of the simulations just 10 of the embryos collided among them at around 3 Myr, 10.5 Myr and 40 Myr. Thus, in the majority of the cases, the resulting planets achieved their final masses at early stages (within 300,000 yr).
@@
As an example of a representative case, in Figure 5 we show the evolution of the planetary mass, semi-major axis and eccentricity of the sample of embryos along the integration time of one of the simulations, together with the final planetary bulk compositions in terms of amount of water in mass. We can see that the planets achieved their final masses and semi-major axis early in their evolution. Once the outermost planets reached its isolation mass, the rest of the planets stopped accreting pebbles and achieve their final configurations. The final masses are between 0.7 and 3 M⊕ and the semi-major axis between 0.01 and 0.1 au. Their eccentricities decreased along the integration time, being less than 0.02 at the end of the simulation. Once the gas disk dissipated, we can see tiny oscillations in semi-major axis and changes up to one order of magnitude in eccentricity. We expect that if we extend the simulations to a Gyr timescale, the orbits of the planets will be completely circularize due to star-planet tidal interactions. The final planetary composition is always rock+H2O. The innermost planet accreted the major amount of water due to the fact that it received a greater amount of embryo impacts which accreted the majority of the pebbles from outside the snowline, while the rest of the planets received fewer embryo impacts and accreted almost half of the pebbles from inside the snowline and close to the other half from outside the snowline.

### v3

@@ 2.2 Dust disk model
Here is the headwind pre-factor that measures the disk pressure gradient (Nakagawa et al., 1986) with , and the gradients of the gas disk surface density and temperature and the gas aspect ratio. In order to take into account the ice sublimation of pebbles when they cross the snowline , we multiply by a factor of 0.5 when . Moreover, we considered that pebbles inside the snowline are dry, while outside the snowline have 50 of water in mass.
The is defined as
@@
We assumed an initial time for our disk model of 1 Myr. We followed the gas disk model proposed by Ida et al. (2016) based on two dominant heating mechanisms: viscous dissipation for the inner disk and stellar irradiation for the outer disk (see Appendix A for a detail description of the model and the parametrization of , , and ).
@@
Unlike Lambrechts & Johansen (2014), who proposed a pebble flux that varies in time depending on the location of the pebble front (radial distance where the pebbles are originated), we assumed a constant pebble flux in time, that varies with the stellar mass. For a star of 0.1 M⊙ we set a pebble flux of M⊕ yr-1, while for a star of 0.3 M⊙ and a star of 0.5 M⊙ we set a pebble flux a factor 6, and a factor 10 higher, respectively, motivated by the fact that the pebble flux scales with the gas density, as suggested in Lambrechts & Johansen (2014). This scale is also consistent with the dust mass and stellar mass relations found by Ansdell et al. (2017). The previous section has demonstrated that these values in compact dust disks correspond to a millimeter dust flux similar to observed values in the unresolved compact disks in Lupus. We note that such values are also in agreement with the order of magnitude of the pebble fluxes estimated by Mulders et al. (2021) for low-mass stars at 1 Myr, and the value proposed recently by Pan et al. (2024) that studied planet formation around 0.1 M⊙ stars. This pebble flux is also consistent with ALMA observations, as demonstrated in Figure 2.
@@
Figure 3 illustrates the pebble model that we used, showing the evolution of the pebble surface density, the pebble aspect ratio and the Stokes numbers at 1 Myr, in which we assumed a star of 0.1 M⊙, a turbulent parameter and the pebble flux of F M. The different slopes of the curves are related to the location of the snowline and to the radial distance that separates the two heating mechanisms proposed in the gas disk model (see Appendix A). We clarify that we assumed the same value for the turbulent parameter as for the gas disk viscosity, in agreement with previous works (e.g., Izidoro et al., 2017; Raymond et al., 2018; Sánchez et al., 2022).
@@
By considering the pebble fluxes mention above, we corroborated that a planet initially located at a few au reach its isolation mass in less than 200,000 yr for a star of 0.1 M⊙ (see Figure 18 as an example), and in less than 100,000 yr for a star of 0.3 M⊙ or 0.5 M⊙, as they are also exposed to a fast inward migration in a gas disk with a low viscosity (see Appendix A for a description of the gas model, Appendix B for planet-disk interactions and Appendix C for the test simulations made to validate the pebble accretion model used). We note that if we calculate the pebble mass available during this period of time from the pebble flux that we proposed, we obtain lower values than the total dust mass available in a disk of 20 au (see Section 2.1). Thus, the pebble flux value proposed is a conservative value which is lower than the maximum value needed for all the dust disk mass within 20 au to be available for the rocky planet assembly (see Section 6 for further discussion).
@@ 3 Planet formation model
low-mass planet-disk interactions, following the gas disk evolution model from Ida et al. (2016), the torques prescriptions from Paardekooper et al. (2010, 2011), and the planetary acceleration corrections prescriptions proposed by Ida et al. (2020), based on purely dynamical friction (Sánchez et al., 2022, see Appendix A and B.1 for a detailed description);
star-planet tidal interactions based on the tidal equilibrium model proposed by Hut (1981), and lately adapted by Bolmont et al. (2011) for brown dwarfs and M dwarfs, which includes the evolution of the stellar radius (Baraffe et al., 2015), and rotational period (Bolmont et al., 2011) (Sánchez et al., 2020, 2022, see Appendix B.2 for a detailed description);
@@
the evolution of the stellar luminosity, and radius based on the stellar evolution models from Baraffe et al. (2015), and the evolution of the rotational period of a star of 0.1, 0.3 and 0.5 M⊙ based on the observations of rotational periods of M dwarfs listed in Bolmont et al. (2011), and in Scholz et al. (2018), for the gas disk model, and the star-planet interactions (see Appendix A and B.2).
@@ 3.1 Formation scenarios
We explored different formation scenarios to study rocky planet formation. First, for our standard model we considered a star of 0.1 M⊙ and a sample of embryos with 0.01 M⊕ initially located beyond the snowline that grow by pebble accretion in a gas disk with low viscosity (, and a disk lifetime of 10 Myr (see the dust disk model in Section 2.2, and the gas model in Appendix A).
@@
as the gas and dust densities are lower in a disk with higher viscosity), to analyze the impact in the resulting planetary masses, and in the migration and orbital decay of the planets. We also assumed a different initial embryo distribution to see the impact in the final fraction of water in planetary masses. Moreover, we let a sample of embryos grow by planetesimal accretion instead of pebble accretion, assuming that the total mass in planetesimalsembryos is equal to the pebble mass available in the disk for accretion (see following sections). Furthermore, we considered higher stellar masses to see how rocky planet formation is scaling with the stellar mass. We considered a star of 0.3 M⊙ and a star 0.5 M⊙. In both cases we studied planet formation by pebble accretion and also by planetesimal accretion in a gas disk with low viscosity () assuming a sample of embryos initially located outside the snowline. For each scenario proposed, we run 10 N-body simulations.
@@ 3.2 N-body code characterization
To run the Mercury code, we chosed the hybrid integrator, which uses a second-order symplectic algorithm to treat interactions between objects with separations greater than 3 RHill, and the Bulirsch-Stöer method to solve close encounters. We run our simulations for 50 Myr each, in order to study the dynamical stability of the system after the gas disk dissipated at 10 Myr. We adopted an integration time step of 0.1 days, in order to be able to map at least ten times the orbit of a planet with an orbital period of 1 day. Furthermore, we consider that an embryo is ejected from the system when it reached a distance of 100 au, and we considered that an embryo had collided with the central star when it is closer than 0.005 au, which corresponds, approximately, to the maximum radius of the host star. We fixed this value for the entire simulation in order to avoid any numerical error for small-periastron orbits.
@@ 3.3 Initial embryo distribution
In the scenarios in which we considered pebble accretion, and an initial embryo distribution outside the snowline, we distributed a total amount of 0.25 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.01 M⊕. The selection of initial embryos with lunar mass is in agreement with several works that consider planetary growth by pebble accretion (e.g., Liu et al., 2019; Coleman et al., 2019) and proposed initially by Kokubo & Ida (1998). Embryos are located beyond the snowline: between 0.5 and 2 au for a central star 0.1 M⊙ ( au), between 1 and 3 au for a central star of 0.3 M⊙ ( au), and between 1.5 and 3.5 au for a central star of 0.5 M⊙ ( au). In each scenario, each seed had a random separation between 5 and 15 Hill Radius, according to their classical isolation mass (e.g., Kokubo & Ida, 2000; Weiss et al., 2018). On the other hand, for the scenario in which we considered embryos located both inside and outside the snowline for a central star of 0.1 M⊙, we distributed 0.5 M⊕ in 50 seeds of protoplanetary embryos, each of them with a mass of 0.01 M⊕. In this case they are located between 0.02 and 1.5 au, randomly distributed between 10 and 15 Hill Radius.
In the scenarios in which we considered planetesimal accretion, we assumed bigger embryos close to the mass of Mars, as in previous works (e.g., Ogihara & Ida, 2009; Coleman et al., 2019). For a central star of 0.1 M⊙, we distributed a total amount of 2.5 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.1 M⊕, located beyond the snowline: between 0.5 and 2 au. For a central star of 0.3 M⊙ and a central star of 0.5 M⊙, we distributed a total amount of 5 M⊕ in 25 seeds of protoplanetary embryos, each of them with a mass of 0.2 M⊕, located between 1 and 3 au and between 1.5 and 3.5 au, respectively. In each simulation, we considered that the embryos initially located inside the snowline were dry, and the ones located outside the snowline present 50 of water in mass
In every case, we considered that all embryos initially have , and . Those values were taken from uniform distributions. The initial values for the rest of the orbital elements, argument of periastron , longitude of the ascending node , and mean anomaly , were randomly determined also from uniform distributions but between 0 and 360 degrees. The physical density of each embryo was fixed to 5 inside the snowline and 1.3 outside the snowline.
@@ 3.4.1 Pebble accretion
For the sample of planetary embryos immerse in a gas disk that grow by pebble accretion, we followed the pebble model proposed by Lambrechts & Johansen (2014). Planetary embryos with different masses accrete pebbles in different regimes. By comparing the pebble accretion radius , and the pebble disk scale height , pebble accretion could be classified into two regimes: the 2D and the 3D regime. If a planetary embryo with mass Mp is large enough (), then it would accrete from the complete layer of pebbles, in the 2D regime; while if the embryo is not large enough (), it would accrete only a fraction of the pebble layer, in the 3D regime. The with the relative velocity between the pebble and the embryo.
The pebble accretion rate is defined as , where, represents the efficiency of pebble accretion (Guillot et al., 2014; Lambrechts & Johansen, 2014). Considering an embryo with an eccentric orbit and an inclination , we calculate such efficiencies following Liu & Ormel (2018) and Ormel & Liu (2018). Thus, the pebble accretion rates in the 2D regime , and in the 3D regime are defined by:
@@
and
are the pebble efficiencies in the 2D and 3D regime, respectively, with and fitting constant of the model, , , and the relative velocity between the embryo and the pebble , with the relative velocity considering an embryo in a circular orbit and the relative velocity considering an embryo in an eccentric orbit, defined as
with the eccentricity of the embryo’s orbit, a fitting constant of the model, the transition mass between the Bondi regime (low impact parameter between the embryo and the pebble) and the Hill regime (high impact parameter) (e.g., Ormel, 2017), relative velocity for a pebble inside Bondi regime, the relative velocity for a pebble inside the Hill regime, and , the transition velocity between the Bondi, and the Hill regimes.
@@
where represents the pebble flux accreted by all the embryos exterior to the radial location of the planet. Each embryo will accrete pebbles until they reach the pebble isolation mass. This mass is considered as the necessary mass that an embryo needs to be able to perturb the local disk structure, forming a pressure bump in the disk exterior to its orbit, stopping the inward migration of pebbles, and limiting the amount of pebbles for embryos orbiting in inner orbits (Lambrechts et al., 2014). The isolation pebble mass can be defined as
@@
As we are working with a sample of embryos, we clarify that each embryo of the simulation will accrete pebbles until it reaches the isolation mass or the total pebble flux available at the location of the planet is zero. If a planet reaches the isolation mass, then we assume that no more pebbles can migrate inward, and thus no other inner planets can continue accreting pebbles (e.g., Lambrechts et al., 2019; Mulders et al., 2021).
@@ 3.4.2 Planetesimal accretion
In the scenarios in which we considered planetesimal accretion, we let the sample of embryos interacts with a sample of 1000 planetesimals. For a star of 0.1 M⊙, each of them has a mass of 0.01 M⊕, while for a star of 0.3 M⊙ the mass of each of them is 0.02 M⊕, and for a star of 0.5 M⊙ is 0.05 M⊕. Their semi-major axis are initially distributed following a power law from the inner edge of the disk until the location where the total amount between embryos, and planetesimals equals the pebble mass available in the disk at 1 Myr, which is between 0.02 au and 2 au for a star of 0.1 M⊙, between 0.065 au and 5 au for a star of 0.3 M⊙ and between 0.075 au and 6 au for a star of 0.5 M⊙.
In every case, the rest of the orbital elements: , , , , and , were estimated from equivalent distributions as for the sample of embryos. The physical density of the planetesimals was fixed to 3 inside the snowline of the system, and 1.5 beyond the snowline. We consider that planetesimales inside the snowline are dry, while outside the snowline present 50 of water in mass.
@@ 4 Simulated systems: Accretion history and dynamical evolution
We present the differences in the resulting simulated planetary systems of each formation scenario (see Table LABEL:tab:table2) in terms of their dynamical evolution and accretion history. Moreover, we compare their final configurations (semi-major axis and eccentricity), masses, amount of water in mass, and multiplicity (see Table 3).
@@ 4.1 Standard model: PA-M01-low-icy
In our standard model we considered a central star of M M⊙, an initial embryo distribution outside the snowline that grow by pebble accretion and a gas disk with low viscosity (). In this scenario, the sample of embryos suffered an efficient inward migration in all the simulations along the disk lifetime. The resulting planets ended up located close-in to the star ( au) in less than 300,000 yr. During this period of time they suffered collisions among embryos and accreted pebbles from outside the snowline for the first 100,000 yr and then from inside the snowline (r au). Once the outermost planet reached the isolation mass, the pebbles stopped migrating inward and, in most cases, the resulting planets achieved their final masses. The resulting planets stopped migrating inward and reminded in compact configurations close to the inner edge of the disk, in quasi-circular and co-planar orbits. In 60 of the simulations, either 1 or 2 collisions with the central star occurred within 500,000 yr. In the rest of the simulations, no embryo collided with the central star. On the other hand, in all the simulations between the 60 and the 88 of the embryos collided among them. In 70 of the simulations, all the collisions among embryos occurred within the first 300,000 yr (while the embryos were also accreting pebbles). In the rest of the simulations just 10 of the embryos collided among them at around 3 Myr, 10.5 Myr and 40 Myr. Thus, in the majority of the cases, the resulting planets achieved their final masses at early stages (within 300,000 yr).
@@
As an example of a representative case, in Figure 5 we show the evolution of the planetary mass, semi-major axis, and eccentricity of the sample of embryos along the integration time of one of the simulations, together with the final planetary bulk compositions in terms of amount of water in mass. We can see that the planets achieved their final masses and semi-major axis early in their evolution. Once the outermost planets reached its isolation mass, the rest of the planets stopped accreting pebbles and achieve their final configurations. The final masses are between 0.7 and 3 M⊕, and the semi-major axis between 0.01 and 0.1 au. Their eccentricities decreased along the integration time, being less than 0.02 at the end of the simulation. Once the gas disk dissipated, we can see tiny oscillations in semi-major axis and changes up to one order of magnitude in eccentricity. We expect that if we extend the simulations to a Gyr timescale, the orbits of the planets will be completely circularize due to star-planet tidal interactions. The final planetary composition is always rock+H2O. The innermost planet accreted the major amount of water due to the fact that it received a greater amount of embryo impacts which accreted the majority of the pebbles from outside the snowline, while the rest of the planets received fewer embryo impacts and accreted almost half of the pebbles from inside the snowline and close to the other half from outside the snowline.